Fermi National Laboratory
 

Accelerator Update for December 18, 2006 - December 20, 2006

 

Monday December 18

The day shift began with Operations monitoring store 5134 (35.32E30) and a Recycler stash, with the Antiproton Source (Pbar) stacking (29.0E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE and NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

At 10:31 AM, experts began repairs on a Booster RF station (BRF6). It was back on line 18 minutes later.

A Preaccelerator expert took down the H- Source for repairs at 10:53 AM. The I- Source has been supplying the accelerators with protons, so this didn’t affect the beam.

The Recycler completed the second and last stash of antiprotons for the day at 10:56 AM.

Operations terminated store 5134 at 12:07 PM.

CDF began a 45-minute period of quiet time at 12:10 PM.

Operations put the Tevatron (TeV) into shot setup at 1:05 PM.

Operations established store 5136 at 3:51 PM with an initial luminosity of 217.0E30.

At 4:29 PM, NuMI reported that their downstream BPMs weren’t working. A NuMI expert investigated and reported that the BPM system was down due to studies. Those studies ended at 5:12 PM and the BPMs were returned to normal operations.

At 9:46 PM, the NuMI horn tripped off on a ground fault indication. It tripped off again while operators were trying to turn it back on. The third try was the charm; the horn reset and turned on.

The Recycler completed its third and last stash for the evening shift at 10:09 PM.

At 10:12 PM, Operations reported that the Booster West Anode Power Supply (WAPS) had tripped off for the first of five times in five minutes. It reset each time.

Operations reported at 11:50 PM that the I- Source reference voltage had tripped off 46 times during the evening shift.

Tuesday December 19

The midnight shift began with Operations monitoring store 5136 (84.07E10) and a stash (146.9E10), with Pbar stacking (49.3E10), and with NuMI and MiniBooNE taking beam.

At 12:47 AM, a Linac RF station (LRF3) tripped off due to a power amplifier crowbar for the first of four times in 11 minutes. It reset each time.

The Recycler completed its third and last stash for the midnight shift at 6:12 AM.

At 6:25 AM, turned off MiniBooNE per schedule. This was done to begin SciBooNE work in MI-12.

At 6:44 AM, Operations halted NuMI beam for about five minutes to perform a crate restore for MI-12.

At 7:55 AM, Operations reported that the I- Source reference voltage had been tripped all shift, but the last trip would not reset. Operations contacted an expert who had the beam back a minute later.

At 9:50 AM, Operations reported that the first Recycler transfer from Pbar was lost. It wasn’t kicked out of MI at the proper time. Operators think it was a sequencer problem.

At 9:57 AM, the MI beam permit dropped due to a MI-8 collimator thermocouple for the first of four times in about 25 minutes. Operations first reported that there were studies going on and then that the trips weren’t real, just motor noise.

Operations reported that power was restored to MiniBooNE. They had a beam permit by 11:55 AM.

At noon, Operations shut off Pbar and prepared the area for a scheduled access. Booster was also prepared for access.

At 12:07 PM, Operations reported that the MiniBooNE horn would not come up to its normal operating current. Operations contacted an expert.

At 12:42 PM, a Preaccelerator expert access the I- Source.

At 12:51 PM, a Linac expert reported that they were starting work on the Linac Debuncher water skid.

At 1:07 PM, Operators tried to turn on tripped vacuum pumps in MI-10, but they wouldn’t stay on.

At 1:23 PM, two NuMI experts were granted a special access into the NuMI Pretarget area.

Pbar came out of access at 1:57 PM.

Operations reported at 2:33 PM that the Pbar rings were on and its permit was green.

Linac experts completed their work on the Linac Debuncher water skid at 2:46 PM.

Operations had beam established to MiniBooNE, NuMI, and Pbar by 2:51 PM.

At 3:50 PM, every TeV RF station came into alarm. Operations contacted a RF expert.

A TeV expert began a study at 5:54 PM.

Operations reported at 7:49 PM that the Pbar Debuncher had been experiencing cooling problems, which had affected the stack rate. Experts had installed a new temperature controller for the Debuncher notch filter oven and had been tuning to compensate for the new device all afternoon.

The evening shift completed its second and last Recycler stash at 9:24 PM.

Operations terminated store 5136 at 10:14 PM.

Operations put the TeV into shot setup at 10:26 PM.

At 11:30 PM, Operations reported that the I- Source had tripped off for the last of many, many times during the course of the shift. It reset.

Wednesday December 20

The midnight shift began with the TeV in shot setup, Operations monitoring the stash, and with MiniBooNE, and NuMI taking beam.

Operations established store 5137 at 12:41 AM with an initial luminosity of 224.95E30.

At 2:59 AM, the cryo system technicians from CHL reported a problem with the wet engine at the TeV F1 sector. Operators investigated and talked with the Cryo Coordinator. The wet engine was operating, but poorly.

The midnight shift completed its last Recycler stash of four at 7:07 AM.

At 7:10 AM, Operations reported that the I- Source had tripped off for the last of many, many, many times during the course of the shift. It reset.

The Plans for Wednesday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/20/2006    email Fermilab